cseltzer:

An unopened cereal box contains `15/28` cubic foot of cereal. If the box is full, and the length and height of the box are `5/7` and `1(1)/(2)` feet, respectively, what is the width of the box? A. `1/2` foot B. `1/7` foot C. `3/5` foot D. `4/7` foot please help

cseltzer:

hello AZ

cseltzer:

can u help me please

AZ:

You should know that Volume = length * width * height They tell you that an unopened cereal box contains 15/28 cubic feet of cereal That is the volume And they tell you that the length is 5/7 feet And that the height is 1 1/2 or 3/2 feet And we want to find out the width

AZ:

So put the numbers into our formula Volume = length * width * height 15/28 = 5/7 * width * 3/2 can you solve for width?

cseltzer:

yes thx

AZ:

What do you get? You can first multiply the fractions 5/7 and 3/2 What is 5/7 * 3/2

AZ:

Remember that \(\dfrac{a}{b} \times \dfrac{c}{d} = \dfrac{a\times c}{b \times d}\)

cseltzer:

?

AZ:

We have 15/28 = 5/7 * 3/2 * width we want to solve for width, right? So first let's simplify it by multiplying the two fractions on the right

cseltzer:

so u mean multiply 3/2 and the width

cseltzer:

hello?

AZ:

No First multiply the length and the height the length is 5/7 and the height is 3/2 What is 5/7 * 3/2

AZ:

We then would divide that from the volume to get the width

cseltzer:

1.07142857143

cseltzer:

that's what i got

AZ:

Don't use a calculator Just multiply the top part of the fraction and the bottom what is 5 * 3 = what is 7 * 2 =

cseltzer:

15 and 14

AZ:

Good so that means \(\dfrac{5}{7} \times \dfrac{3}{2} = \dfrac{5\times 3}{7\times 2} = \dfrac{15}{14}\) Does that make sense?

AZ:

So now we have Volume = length * width * height 15/28 = 5/7 * 3/2 * width and we said that 5/7 * 3/2 = 15/14 so we get 15/28 = 15/14 * width

AZ:

Are you following along still? So what does the width have to be if we're going to multiply 15/14 to make it 15/28 What are we multiply the 14 in the denominator by to get 28?